OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. A. y = -x + 6 B. y = x - 6 C. y = -x - 2 D. y = x + 2

OpenStudy (anonymous):

@KamiBug @shaadykhamlichi

OpenStudy (shaadykhamlichi):

D

OpenStudy (shaadykhamlichi):

trust me it's D

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. A. y - 1 = 4(x - 3) B. y + 1 = 4(x + 3) C. y - 3 = 4(x + 1) D. y + 3 = 4(x - 1)

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

@KamiBug @felavin @MakaylaTracy @panchajanya @Coolsector

OpenStudy (kamibug):

B. y + 1 = 4(x + 3)

OpenStudy (anonymous):

use point slope form

OpenStudy (anonymous):

ok thanks :*

OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) - 5? A. The graph shifts 5 units up. B. The graph shifts 5 units down. C. The graph shifts 5 units left. D. The graph shifts 5 units right.

OpenStudy (anonymous):

y-y1 = m(x-x1) here m is 4and (x1.y1) is (-3,-1)

OpenStudy (anonymous):

C right ?

OpenStudy (kamibug):

B. The graph shifts 5 units down.

OpenStudy (anonymous):

okok

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y =1/4x + 2. A. y = 4x - 1 B. y = 4x - 3 C. y = 1/4x - 3 D. y = 1/4x - 1

OpenStudy (kamibug):

D. y = 1/4x - 1

OpenStudy (anonymous):

thank you ;)

OpenStudy (kamibug):

Sure :)